OpenStudy (anonymous):

square root of -5 multiplied by square root of 5

OpenStudy (j2lie):

-5 and 5 does not have a perfect square root.

OpenStudy (anonymous):

\[\sqrt{-5} \times \sqrt{5}\]

OpenStudy (anonymous):

So what does that mean?

OpenStudy (j2lie):

The square root is between 2 and 3.

OpenStudy (anonymous):

I dont get it

OpenStudy (j2lie):

the square root is either 2.5 or below.

OpenStudy (kropot72):

\[\large \sqrt{-1}=i\] The square root of negative one is the imaginary number i. Factorising the square root of negative 5 we get: \[\large \sqrt{-5}\times \sqrt{5}=(\sqrt{5}\times \sqrt{5})i\]

OpenStudy (kropot72):

The get the answer, which is an imaginary number, just simplify the bracketed terms.

OpenStudy (anonymous):

How do I simplify bracketed terms

OpenStudy (kropot72):

Multiply the square root of 5 by the square root of 5.

OpenStudy (teddyiswatshecallsme):

I think its. \[\sqrt{-5} {\times} \sqrt{5} = \sqrt[4]{-25}\]

OpenStudy (j2lie):

That makes more sense. @Teddyiswatshecallsme

OpenStudy (anonymous):

Im confused now.

OpenStudy (kropot72):

\[\large (\sqrt{5}\times \sqrt{5})i=5i\]

OpenStudy (anonymous):

Its -5

OpenStudy (kropot72):

The correct result must be an imaginary number.

OpenStudy (anonymous):

So is @Teddyiswatshecallsme right or @kropot72

OpenStudy (kropot72):

Has anybody given Teddy a medal?

OpenStudy (anonymous):

No

OpenStudy (kropot72):

@basketball305 Who will judge which is the correct answer?

OpenStudy (zzr0ck3r):

@kropot72 is absolutely correct. Trust the person that shows there work, because if they are wrong you can find it in their work \(\sqrt{5}\sqrt{-5}=\sqrt{5}\sqrt{-1*5}=\sqrt{5}\sqrt{5}\sqrt{-1}=5\sqrt{-1}=5i\)

OpenStudy (zzr0ck3r):

their*

OpenStudy (anonymous):